Grade 11General Physics

What is the angle between two vector forces of equal magnitude such that the resultant is one third as much as either of the original forcse?

Praveen
10 Years agoGrade 11

1 Answer

Vikas TU
10 Years ago
Hi Praveen,Let the force vectors be A and B.Then |A| =|B| ….......................(1) {Given}And also,|A +B| = |A|Squaring both sides,|A|^2 + |B|^2 + 2.A.B = |A|^2|B|^2 + 2A.B = 0..........................(2)Now Angle between the tqo vectors is:cos@ = A.B/|A|.|B| = A.B/|B|^2 (Given = |A| =|B|). From eqn. (2)A.B/|B|^2 = -1/2 = cos@@ = 120 degree.